Mudvayne, Inc. is trying to determine its cost of debt. The firm has a debt issue outstanding with 18 years to maturity that is quoted at 107 percent of face value. The issue makes semiannual payments and has an embedded cost of 6 percent annually. What is the company's pretax cost of debt? If the tax rate is 35 percent, what is the aftertax cost of debt?

Excel Sheet:

A B

1 Settlement 1/1/00

2 Maturity 1/1/18

3 Price (% of Par) 107

4 Coupon rate 6%

5 Payments per year 2

6 Tax rate 35%

Do not hardcode values, need formulas. (I've already tried this formula =RATE(36,30,-1070,1000,0), it doesn't work. It should work, but it doesn't. I don't think the system likes how I've put in the formula, maybe it's wanting the formula a specific way).

Pretax Cost - ?

Aftertax cost of debt - ?
Answer #1

Numer of years = 18

Number of Semi Annual Periods = 18 * 2 = 36

Semi Annual Coupon Payment = 6% * 1,000 * 0.5 = 30

FV = 1,000

PV = 1,070

Book1 - Excel Insert Page LayoutFormulas Data Review View Tell me what you want to d.o.. File Home Cut Percentage Calibri For

As you can see, by putting the values in excel,

we get I = 2.69%

YTM = 2.69% * 2

YTM = 5.39%

Pre-tax cost of debt = 5.39%

After tax cost of debt = 5.39% * (1 - 35%)

After tax cost of debt = 3.50%